OpenStudy (anonymous):

Multiply and simplify. 6 1/4*4/5 A.1/5 B.5 C.1/4 D.4

OpenStudy (anonymous):

@Agl202

OpenStudy (mathwizzard3):

I can help

OpenStudy (agl202):

25/4 × 4/5 = (25 × 4) / (4 × 5) = 100/20 = 5/1 = B. 5